OpenStudy (anonymous):

3(2x + 5) - x = [ ]x + [ ] what goes into the box to make it an infinite number of solutions.

Vocaloid (vocaloid):

hint: distribute the 3 on the left side

Vocaloid (vocaloid):

3*2x = ?

OpenStudy (anonymous):

6x?

Vocaloid (vocaloid):

yeah now 3*5 = ?

OpenStudy (anonymous):

15

OpenStudy (anonymous):

thx!!!!!!!!!!!!!!

Vocaloid (vocaloid):

good now we have 6x + 15 - x now combine like terms 6x - x = ?

OpenStudy (anonymous):

still wait 6?

Vocaloid (vocaloid):

not quite Start 6x take away one x how many x's are left?

OpenStudy (anonymous):

0 now.. sooo. 6

Vocaloid (vocaloid):

nope not quite...

Vocaloid (vocaloid):

let me put it a different way

Vocaloid (vocaloid):

start with 6 dogs take away one dog how many dogs are left?

OpenStudy (anonymous):

5.... oh wait sooo........... 5 or 5x?!?!

Vocaloid (vocaloid):

5x

OpenStudy (anonymous):

OKAY...

Vocaloid (vocaloid):

6x - x = 5x so now our final answer is 5x - 15

Vocaloid (vocaloid):

sorry

Vocaloid (vocaloid):

5x + 15

OpenStudy (anonymous):

YAY thx!!!!